OpenStudy (anonymous):

very desperate! find the sum of the first 50 terms of the sequence an=3n+2

myininaya (myininaya):

\[1+2+\cdots+k=\frac{k(k+1)}{2} \\ \text{ multiplying both sides by a constant } c \text{ gives } \\ c(1+2+\cdots+ k)=c \cdot \frac{k(k+1)}{2} \\ \\ \text{ also if you are adding an } k \text{ amount of two's together } \\ \text{ like } 2+2+2+ \cdots +2 =2k\]

myininaya (myininaya):

so we can look at this seperate you have \[a_n=b_n+c_n \\ \sum a_n=\sum(b_n+c_n)=\sum b_n +\sum c_n\] that big looking weird symbol just means the sum of the terms \[\text{ anyways } 1+2+3 +\cdots +50=?\]

OpenStudy (anonymous):

how do i plug the sequence into the sigma?

OpenStudy (anonymous):

or the equation?

myininaya (myininaya):

\[\sum_{n=1}^{50}(3n+2)=\sum_{n=1}^{50} (3n)+\sum_{n=1}^{50}2 =3 \sum_{n=1}^{50} n +\sum_{n=1}^{50} 2\] so this is why I'm asking you to first calculate 1+2+3+...+50 I have given you a formula above to this

myininaya (myininaya):

this notation I wrote \[\sum_{n=1}^{50}(3n+2)=\sum_{n=1}^{50} (3n)+\sum_{n=1}^{50}2 =3 \sum_{n=1}^{50} n +\sum_{n=1}^{50} 2 \\ \text{ can be written as } \\ 3(1+2+3 + \cdots +50)+(2 +2 + \cdots +2)\] that is suppose to be fifty amount of twos

OpenStudy (anonymous):

is it 3925?

myininaya (myininaya):

that is right \[3 \cdot \frac{50(51)}{2}+2(50) \text{ using both of the formulas I mentioned \in the first post }\]

OpenStudy (anonymous):

oh okay! so that's the answer t the question?

myininaya (myininaya):

what I mean by " that is right" is "that is correct"

OpenStudy (anonymous):

thank you!